cookiemonster1:

Select all the points that are on the graph of the line 2x + 4y = 20. (0,5) (0,10) (1,2) (1,4) (5,0) (10,0)

ramen:

For the easy way, why don't you try graphing this eqation?

ramen:

And then see what points lie on this line.

cookiemonster1:

i dont have paper though

ThisGirlPretty:

Have you tried this site? https://www.desmos.com/calculator

ramen:

Well if you don't have paper, then I could teach you the mathematical way.

ramen:

If we plug in a coordinate into the equation 2x + 4y = 20, we know if the point lies on the line if the right and left side equal each other, for example 20=20.

ramen:

For (0,5) we would plug in 0 into the x, so 2(0) and 5 into the y so 4(5). The equation now is 2(0) + 4(5)= 20.

ramen:

Well since 2(0) + 4(5) = 20, this coordinate lies on the line. You can do this for each of the coordinates; x value into x and y value into y.

ramen:

hope this helps

cookiemonster1:

thanks

ramen:

no problem